# Building Gradient by Gradient: Decentralised Energy Functions for Bimanual Robot Assembly

### v1

@@ Abstract
There are many challenges in bimanual assembly, including high-level sequencing, multi-robot coordination, and low-level, contact-rich operations such as component mating. Task and motion planning (TAMP) methods, while effective in this domain, may be prohibitively slow to converge when adapting to disturbances that require new task sequencing and optimisation. These events are common during tight-tolerance assembly, where difficult-to-model dynamics such as friction or deformation require rapid replanning and reattempts. Moreover, defining explicit task sequences for assembly can be cumbersome, limiting flexibility when task replanning is required. To simplify this planning, we introduce a decentralised gradient-based framework that uses a piecewise continuous energy function through the automatic composition of adaptive potential functions. This approach generates sub-goals using only myopic optimisation, rather than long-horizon planning. It demonstrates effectiveness at solving long-horizon tasks due to the structure and adaptivity of the energy function. We show that our approach scales to physical bimanual assembly tasks for constructing tight-tolerance assemblies. In these experiments, we discover that our gradient-based rapid replanning framework generates automatic retries, coordinated motions and autonomous handovers in an emergent fashion.
@@ I Introduction
Figure 1: A decentralised gradient-based planning approach for real-world bimanual assembly. Each robot creates a piecewise continuous energy function through the automatic switching between myopic functions. The planner facilitates rapid replanning, where reattempts, coordination between arms and human/robot collaboration emerge naturally. A summary of the alogrithm and results is found at https://youtu.be/0GIxA7_Ud2k.
@@
We address these challenges by introducing a decentralised gradient-based planner for bimanual assembly, where sub-tasks arise automatically as we optimise the program. Discrete task-specific elements of our planning program are projected into a piecewise continuous energy function. A single energy function is created by selecting from a set of active potential functions. We define an energy function for each robot arm to every component in the assembly. At every planning step, each robot selects an energy function with the smallest non-zero magnitude. This energy function corresponds to the component which is closest to its desired location, but is not in the goal pose. Once a part is in the correct location, the potential field adjusts, and the robot moves onto a new sub-goal automatically. Sub-goals are generated at run time without a task planner or sketch.
@@ II Related Works
The de facto approach for bimanual assembly planning is task and motion planning (TAMP) [1, 4, 15], which combines discrete task planning (e.g., contact timings and scheduling) with continuous trajectory optimisation. However, these methods rely on handcrafted domain knowledge and therefore struggle with unmodelled dynamics seen in the real-world. This is common in tight-tolerance assembly, where effects such as friction or deformation can disrupt execution, forcing robots to reattempt or resequence [5]. Toussaint et al. [16] addresses replanning limitations by integrating the discrete task planning components into a singular continuous program. This approach finds trajectories by optimising robot states and timings between task-specific constraints, given a sketch of task constraints. This is a fast elegant TAMP approach and facilitates rapid replanning. However, this method requires a task sketch, which typically cannot be rearranged at planning time due to runtime of MCTS [17] being in the order of seconds rather than milliseconds.
@@ V Evaluation
Our evaluations of our approach are designed to investigate the following:
@@
Figure 4: Our approach optimises both subgoals and motions concurrently using gradient descent for assembly. Starting at the top row, the robot grasps a component in each hand. These are moved to their target locations. The right arm regrasps its component to make slight adjustments to its pose in the top right figure. Once the first two pieces are correctly located, the robot moves its arms to other components for assembly, as shown in the middle row. The bottom row shows the insertion of the last two components. Coordination arises between the arms automatically as one component disturbs another. As the left hand adjusts the final pose of both components, the left arm retracts, as shown in the bottom row, middle figure. A video of our results can be found at the following link: https://youtu.be/0GIxA7_Ud2k.
@@ V-B Convergence
V-C Real-World Bimanual Setup
@@
Figure 5: In the top sub-figure, the robot reacts to a human operator disassembling the assembly. The robot rebuilds to the previous goal accurately enough for the operator to insert a pin into the assembly. In the bottom sub-figure, the robot’s left hand moves a component into the right arm’s workspace. The right arm continues to move this piece into place. An accompanying video demonstrates these results and is found at https://youtu.be/0GIxA7_Ud2k.
V-D Real-World Assembly

### v2

@@ Abstract
There are many challenges in bimanual assembly, including high-level sequencing, multi-robot coordination, and low-level, contact-rich operations such as component mating. Task and motion planning (TAMP) methods, while effective in this domain, may be prohibitively slow to converge when adapting to disturbances that require new task sequencing and optimisation. These events are common during tight-tolerance assembly, where difficult-to-model dynamics such as friction or deformation require rapid replanning and reattempts. Moreover, defining explicit task sequences for assembly can be cumbersome, limiting flexibility when task replanning is required. To simplify this planning, we introduce BGBG, a decentralised gradient-based framework that uses a piecewise continuous energy function through the automatic composition of adaptive potential functions. This approach generates sub-goals using only myopic optimisation, rather than long-horizon planning. It demonstrates effectiveness at solving long-horizon tasks due to the structure and adaptivity of the energy function. We show that our approach scales to physical bimanual assembly tasks for constructing tight-tolerance assemblies. In these experiments, we discover that our gradient-based rapid replanning framework generates automatic retries, coordinated motions and autonomous handovers in an emergent fashion.
@@ I Introduction
Figure 1: A decentralised gradient-based planning approach for real-world bimanual assembly. Each robot creates a piecewise continuous energy function through the automatic switching between myopic functions. The planner facilitates rapid replanning, where reattempts, coordination between arms and human/robot collaboration emerge naturally. A summary of the alogrithm and results is found at https://youtu.be/Kw99FtEhZB8.
@@
We address these challenges by introducing Building Gradient by Gradient (BGBG), a decentralised gradient-based planner for bimanual assembly, where sub-tasks arise automatically as we optimise the program. Discrete task-specific elements of our planning program are projected into a piecewise continuous energy function. A single energy function is created by selecting from a set of active potential functions. We define an energy function for each robot arm to every component in the assembly. At every planning step, each robot selects an energy function with the smallest non-zero magnitude. This energy function corresponds to the component which is closest to its desired location, but is not in the goal pose. Once a part is in the correct location, the potential field adjusts, and the robot moves onto a new sub-goal automatically. Sub-goals are generated at run time without a task planner or sketch.
@@ II Related Works
The de facto approach for bimanual assembly planning is task and motion planning (TAMP) [15, 4, 16], which combines discrete task planning (e.g., contact timings and scheduling) with continuous trajectory optimisation. However, these methods rely on handcrafted domain knowledge and therefore struggle with unmodelled dynamics seen in the real-world. This is common in tight-tolerance assembly, where effects such as friction or deformation can disrupt execution, forcing robots to reattempt or resequence [5]. Toussaint et al. [17] addresses these limitations by integrating the discrete task planning components into a singular continuous program. This approach finds trajectories by optimising robot states and timings between task-specific constraints, given a sketch of task constraints. This is a fast elegant TAMP approach and facilitates rapid replanning. However, this method requires a task sketch, which typically cannot be rearranged at planning time due to runtime of MCTS [18] being in the order of seconds rather than milliseconds. The Multi-Robot Path Planning Benchmark (MRPP) [15] proposes solving both task and motions together. This method jointly explores robot configurations and discrete task-transition modes for pick-and-place tasks in simulation to enable multi-robot, multi-goal planning with grasping and placement actions. While the approach handles complex multi-modal problems, it relies heavily on global sampling and sequential mode transitions, which makes it sensitive to poor task-sequence exploration, leading to long planning times.
@@ V Evaluation
We evaluate BGBG to investigate the following:
@@
    How quickly does BGBG plan tasks and motions per component?
  3. 3.
@@
Figure 4: Our approach optimises both subgoals and motions concurrently using gradient descent for assembly. Starting at the top row, the robot grasps a component in each hand. These are moved to their target locations. The right arm regrasps its component to make slight adjustments to its pose in the top right figure. Once the first two pieces are correctly located, the robot moves its arms to other components for assembly, as shown in the middle row. The bottom row shows the insertion of the last two components. Coordination arises between the arms automatically as one component disturbs another. As the left hand adjusts the final pose of both components, the left arm retracts, as shown in the bottom row, middle figure. A video of our results can be found at the following link: https://youtu.be/Kw99FtEhZB8.
@@ V-B Convergence
V-C Task And Motion Planning Speed Comparison
The task and motion replanning speed of BGBG is evaluated and compared to other methods. Rapid task-level planning is essential for reliable robot assembly, especially when tight tolerances make reattempts common. The methods selected for comparison are Hartmann et al.’s Multi-Robot Path Planning Benchmark (MRPP) [15] and the SAT solver used for task planning in RAMP [5]. MRPP jointly explores robot configurations and discrete task-transition modes to solve for task and motion plans for pick and place problems. The SAT solver in RAMP finds a detailed plan at both the coarse and fine levels for a single robot assembling components under the assumption that one of the components is rigidly fixed. The all experiments, eight components are selected and initialised in random configurations. The initial conditions are the same for all the solvers.
The per-component solver times for BGBG are compared to four solver setups. The first instance is MRPP provided with a task sequence, and a goal configuration. In this situation MRPP operates and multi-robot motion planner and finds trajectories for two arms and for all the components. In this mode, MRPP is operating as a multi-robot pick and place solver and does not plan how to mate the components. The per-component solver times for eight assemblies planning for eight components is reported in Fig. 5 in the first two columns. The BGBG and MRPP (Pick) formulations are both rapid solvers. However, BGBG is in general a faster solver. In our next experiment, we use the full TAMP version of MRPP, which samples insertions for component mating and task sequences. In Fig. 5, the spread of solver durations are reported in the middle column. The additional sampling slows the optimisation durations, and can cause entrapment in local minima.
A comparison between a purely task-based planner and BGBG is executed. This comparison investigates whether the RAMP [5] SAT solver can generate task sequences quickly enough for replanning. The RAMP SAT solver assumes that one component is rigidly fixed to a surface and plans under the assumption of a single manipulator. The planner creates a coarse plan (e.g. where and when should each component be) and a fine-grained plan to organise pre and post-insertion tasks. The durations of both operating modes for the assembly is reported in the right two columns of Fig. 5. This task-solver is significantly slower than both BGBG and MRPP.
Figure 5: Timing comparison for planning assembly motions. BGBG is our method, MRPP represents [15] and SAT is the SAT solver from [5]. The MRPP Pick does not consider component mating, whilst MRPP TAMP version does. Lower planning durations lead to faster task replanning for reattempts during assembly.
Figure 6: In the top sub-figure, the robot reacts to a human operator disassembling the assembly. The robot rebuilds to the previous goal accurately enough for the operator to insert a pin into the assembly. In the bottom sub-figure, the robot’s left hand moves a component into the right arm’s workspace. The right arm continues to move this piece into place. An accompanying video demonstrates these results and is found at https://youtu.be/Kw99FtEhZB8.
V-D Real-World Bimanual Setup
@@
V-E Real-World Assembly